# Classical and quantum chaotic synchronization in coupled dissipative time crystals

### v1

@@ Abstract
We investigate the dynamics of two coherently coupled dissipative time crystals. In the classical mean-field limit of infinite spin length, we identify a regime of chaotic synchronization, marked by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one. At the boundary of this regime, the Pearson coefficient varies abruptly, marking a crossover between staggered and uniform $z$-magnetization. To address finite-size quantum dynamics, we employ a quantum-trajectory approach and study the trajectory-resolved expectations of subsystem $z$-magnetizations. Their histograms over time and trajectory realizations exhibit maxima that undergo a staggered-to-uniform crossover analogous to the classical one. In analogy with the classical case, we interpret this behavior as quantum chaotic synchronization, with dissipative quantum chaos evidenced by the steady-state density matrix exhibiting Gaussian Unitary Ensemble statistics. The locations of the classical and quantum crossover differ, reflecting the noncommutativity of the infinite-spin and infinite-time limits and the impact of entanglement, quantified via the entanglement entropy between subsystems.
@@ 1 Introduction
To probe quantum effects at finite spin magnitude, we employ the quantum-trajectory method and analyze trajectory-resolved $z$-magnetizations. Evaluating for each subsystem the histogram of the $z$-magnetization expectations, as they change over time and trajectory realizations, one can see a maximum that displays a sharp crossover from a staggered to a uniform regime across the two subsystems. This behavior closely echoes the classical case and suggests the emergence of quantum chaotic synchronization. This interpretation is strengthened by the steady-state density matrix, whose spectrum exhibits Gaussian unitary ensemble statistics that marks quantum chaos, and by the behavior of the bipartite entanglement entropy, highlighting the central role of entanglement in quantum synchronization [13, 4, 5, 40]. Notably, the crossover points in the quantum and classical cases do not coincide, reflecting the noncommutativity of the infinite-time and infinite-spin limits: While the classical system exhibits persistent oscillations, the quantum dynamics relaxes to a nonequilibrium steady state [41]. Nevertheless, both regimes display qualitatively similar synchronization crossovers, although the physics is different and in the quantum case a role in the crossover is played by entanglement.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $\Omega/2+\Gamma=2$ and on the right one $\Omega/2+\Gamma=1.5$.
@@
and plot its histograms in Fig. 6(c). All histograms have a maximum, and we plot its position as a function of $\Gamma$ in Fig. 7. We can clearly observe the crossover from the staggered to the uniform magnetized regime in the behavior of $(S_{A}^{z})^{\rm max}$ and $(S_{B}^{z})^{\rm max}$ [Fig. 7(a,c)], also accompanied by a peak in the behavior of the maximum point of the entanglement entropy $\mathcal{E}^{\text{max}}$ [Fig. 7(b,d)].
Figure 6: (a) Histogram of $(S_{A}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (b) Histogram of $(S_{B}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (c) Histogram of the entanglement entropy $\mathcal{E}^{r}(t)$ as it changes along time $t$ and across trajectories $r$. We fix the constraint $\Omega/2+\Gamma=1.5$, $\delta t=10^{-4}$ and $S=20$.
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $\Omega/2+\Gamma=2$ in the left column and $\Omega/2+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$.
@@
We notice that the threshold for this crossover in the quantum case is smaller than in the mean field case [compare Fig. 7(a,c) with the corresponding Fig. 2(b,c)]. The point is that the infinite-$S$ limit and the infinite time-limit do not commute. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [52, 41]). Thus, it is not strange that the crossover occurs at different values of $\Gamma$ for the mean-field infinite-$S$ regime and for the long-time limit of the quantum finite-$S$ case. What is remarkable is that the crossover is present in both scenarios, despite the non-commutativity of the limits and their different long-time behaviors; additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent (LLE), shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 8). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.
@@
Figure 8: Average level spacing ratio of the NESS density matrix versus $\Gamma$ for different values of $S$, compared with the average level spacing ratio for the Gaussian Unitary Ensemble as evaluated in Ref. [51]. The flatness of the curve and the agreement with the GUE value increase for increasing subsystem spin magnitude $S$. We fix the constraint $\Omega/2+\Gamma=2$ and evaluate $\hat{\rho}_{\rm NESS}$ as $\hat{\rho}_{t}$ with $t=100$ in units of $1/\kappa$. In Eq. (21) we average over $N_{\rm r}=48$ quantum trajectories and take $\delta t=10^{-4}$.
@@ 5 Conclusions
We then turned to finite-$S$ systems, where quantum fluctuations become relevant. Using a quantum-trajectory approach, we analyzed subsystem magnetizations along stochastic trajectories and their distributions over time and trajectory realizations. The maxima of these histograms exhibit a sharp crossover between staggered and uniform magnetization, closely resembling the mean-field case and indicating a quantum analogue of chaotic synchronization. Moreover, dissipative quantum chaos is revealed by Gaussian unitary ensemble statistics in the nonequilibrium steady-state density matrix [50]. Beyond that, the entanglement entropy between subsystems displays a correlated feature at the crossover, showing that entanglement contributes to the quantum synchronization process [13, 4, 5].
In summary, this model displays classical chaotic synchronization and an analogous quantum phenomenon, which we can call quantum chaotic synchronization. Interestingly, the crossover between the staggered and the uniform magnetization in the mean-field classical chaotic synchronization occurs at a different point in parameter space than the finite-$S$ threshold between staggered and uniform magnetization in the quantum case. This difference arises because the limits $S\to\infty$ and $t\to\infty$ do not commute in this quantum dynamics. In the mean-field case, the $S\to\infty$ limit is taken first, resulting in persistent oscillations that can be periodic or aperiodic [39]. In the quantum case, the $t\to\infty$ limit comes first and the system always relaxes to a NESS. Remarkably, the NESS shares qualitative features with the mean-field state, despite their fundamental differences and the presence of entanglement only in the quantum case. The noncommutativity of these limits further explains the discrepancies in parameter regimes where classical and quantum chaos appear.

### v2

@@ Abstract
We investigate the dynamics of two coherently coupled dissipative time crystals. In the classical mean-field limit of infinite spin length, we identify a regime of chaotic synchronization, marked by a positive largest Lyapunov exponent and a Pearson correlation coefficient close to one. At the boundary of this regime, the Pearson coefficient varies abruptly, marking a crossover between staggered and uniform $z$-magnetization. To address finite-size quantum dynamics, we employ a quantum-trajectory approach and study the trajectory-resolved expectations of subsystem $z$-magnetizations. Their histograms over time and trajectory realizations exhibit maxima that undergo a staggered-to-uniform crossover analogous to the classical one. In analogy with the classical case, we interpret this behavior as quantum chaotic synchronization, with dissipative quantum chaos evidenced by the steady-state density matrix exhibiting Gaussian Unitary Ensemble statistics. The locations of the classical and quantum crossover are consistent with each other, reflecting quantum-classical correspondence despite the noncommutativity of the infinite-spin and infinite-time limits and the impact of entanglement, quantified via the entanglement entropy between subsystems.
@@ 1 Introduction
To probe quantum effects at finite spin magnitude, we employ the quantum-trajectory method and analyze trajectory-resolved $z$-magnetizations. Evaluating for each subsystem the histogram of the $z$-magnetization expectations, as they change over time and trajectory realizations, one can see a maximum that displays a sharp crossover from a staggered to a uniform regime across the two subsystems. This behavior closely echoes the classical case and suggests the emergence of quantum chaotic synchronization. This interpretation is strengthened by the steady-state density matrix, whose spectrum exhibits Gaussian unitary ensemble statistics that marks quantum chaos, and by the behavior of the bipartite entanglement entropy, highlighting the central role of entanglement in quantum synchronization [13, 4, 5, 40]. Notably, the crossover points in the quantum and classical cases coincide, despite the noncommutativity of the infinite-time and infinite-spin limits: While the classical system exhibits persistent oscillations, the quantum dynamics relaxes to a nonequilibrium steady state [41]. Nevertheless, both regimes display a qualitatively and quantitatively similar synchronization crossover, although the physics is different and in the quantum case a role in the crossover is played by entanglement.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $2\Omega+\Gamma=2$ and on the right one $2\Omega+\Gamma=1.5$.
@@
and plot its histograms in Fig. 6(c). All histograms have a maximum, and we plot its position as a function of $\Gamma$ in Fig. 7. We can clearly observe the crossover from the staggered to the uniform magnetized regime in the behavior of $(S_{A}^{z})^{\rm max}$ and $(S_{B}^{z})^{\rm max}$ [Fig. 7(a,c)], also accompanied by a local minimum in the behavior of the maximum point of the entanglement entropy $\mathcal{E}^{\text{max}}$ [Fig. 7(b,d)].
Figure 6: (a) Histogram of $(S_{A}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (b) Histogram of $(S_{B}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (c) Histogram of the entanglement entropy $\mathcal{E}^{r}(t)$ as it changes along time $t$ and across trajectories $r$. We fix the constraint $2\Omega+\Gamma=1.5$, $\delta t=10^{-4}$ and $S=20$.
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $2\Omega+\Gamma=2$ in the left column and $2\Omega+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$.
Figure 8: Heat map of $s=\operatorname{sign}(\braket{\overline{m_{A}^{z}}}\braket{\overline{m_{B}^{z}}})$ as shown in Fig. 3(f) where the lines $2\Omega+\Gamma=2$ (upper dotted line) and $2\Omega+\Gamma=1.5$ (lower dashed line) are drawn.
We notice that the threshold for this crossover in the quantum case is consistent with the mean field case, as we can see comparing Fig. 7(a,c) with Fig. 8. This figure is the same as Fig. 3(f) and depicts $s=\operatorname{sign}(\braket{\overline{m_{A}^{z}}}\braket{\overline{m_{B}^{z}}})$ that is -1 in case of staggered magnetization and +1 in case of uniform magnetization. On this plot we have drawn the lines $2\Omega+\Gamma=2$ (upper dotted line) and $2\Omega+\Gamma=1.5$ (lower dotted-dashed line). We see that the upper line crosses the classical boundary between staggered and uniform magnetization for $\Gamma\simeq 1.25$ that is consistent with the value where uniform magnetization sets in the corresponding quantum case [see Fig. 7(a)]. The same occurs for the lower line that crosses the boundary at $\Gamma\simeq 1$, consistently with the corresponding quantum case shown in Fig. 7(c). So we can see a classical-quantum correspondence in the threshold between staggered and uniform magnetization.
@@
This happens although the infinite-$S$ limit and the infinite time-limit do not commute. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [52, 41]). So it is remarkable is that the crossover is present in both scenarios, and occurs at the same point, despite the non-commutativity of the limits and their different long-time behaviors. Additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent (LLE), shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 9). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.
@@
Figure 9: Average level spacing ratio of the NESS density matrix versus $\Gamma$ for different values of $S$, compared with the average level spacing ratio for the Gaussian Unitary Ensemble as evaluated in Ref. [51]. The flatness of the curve and the agreement with the GUE value increase for increasing subsystem spin magnitude $S$. We fix the constraint $2\Omega+\Gamma=2$ and evaluate $\hat{\rho}_{\rm NESS}$ as $\hat{\rho}_{t}$ with $t=100$ in units of $1/\kappa$. In Eq. (21) we average over $N_{\rm r}=48$ quantum trajectories and take $\delta t=10^{-4}$.
@@ 5 Conclusions
We then turned to finite-$S$ systems, where quantum fluctuations become relevant. Using a quantum-trajectory approach, we analyzed subsystem magnetizations along stochastic trajectories and their distributions over time and trajectory realizations. The maxima of these histograms exhibit a sharp crossover between staggered and uniform magnetization, closely resembling the mean-field case and indicating a quantum analogue of chaotic synchronization. Moreover, dissipative quantum chaos is revealed by Gaussian unitary ensemble statistics in the nonequilibrium steady-state density matrix [50]. Beyond that, the entanglement entropy between subsystems displays a correlated feature (a local minimum) at the crossover, showing that entanglement contributes to the quantum synchronization process [13, 4, 5].
In summary, this model displays classical chaotic synchronization and an analogous quantum phenomenon, which we can call quantum chaotic synchronization. Interestingly, the crossover point between the staggered and the uniform magnetization in the mean-field classical chaotic synchronization is consistent with the one in the quantum regime, confirming quantum-classical correspondence. This is an interesting finding because the limits $S\to\infty$ and $t\to\infty$ do not commute in this quantum dynamics. In the mean-field case, the $S\to\infty$ limit is taken first, resulting in persistent oscillations that can be periodic or aperiodic [39]. In the quantum case, the $t\to\infty$ limit comes first and the system always relaxes to a NESS. Remarkably, the NESS shares qualitative and quantitative features with the mean-field state, despite their fundamental differences and the presence of entanglement only in the quantum case. The noncommutativity of these limits explains the discrepancies in parameter regimes where classical and quantum chaos appear.